OpenStudy (anonymous):

Product A is a 12-ounce bottle of generic popcorn oil that sells for $1.39. Product B is a 24-ounce popcorn oil bottle that costs $2.79. Which is the better buy?

OpenStudy (galleara):

product A because if you multiply the price by 2, you get $2.78 which saves a whole wopping .01 cent haha

OpenStudy (texaschic101):

Product A : 12/1.39 = 1/x cross multiply 12x = 1.39 x = 1.39/12 x = .115 Product B : 24/2.79 = 1/x cross multiply 24x = 2.79 x = 2.79/24 x = .116 Product A is cheaper.....not by much :)